# Automated System-level Testing of Unmanned Aerial Systems

### v1

@@ 5.6.1 External Validity Threat
The possible threats to the validity of our experiment are discussed individually in the following subsections. The external threat to validity is related to the generalization of the experiment results. To reduce the chance of external validity threat, the experiment was performed using two primary subsystems of UAS, an autopilot (ArduCopter) and the CDS of GCS (GCS-CDS). ArduCopter is a good representative open-source autopilot system that is widely used in research and development such as in [34] and [17]. The ArduCopter supports a wide variety of multi-copter UAVs including a helicopter, tri-copter, quad-copter, hexacopter, etc. Moreover, this is also compatible with a wide range of hardware e.g., Pixhawk, CUAV v5 Plus, OpenPilot Revolution, etc. The GCS-CDS is an industrial case study containing widely used CDS elements. The results of our experiment may not be generalizable for all subsystems of UAS, however, this threat typically exists in every experiment.
@@ 5.6.2 Internal Validity Threat
The parameter tuning of the training algorithm presents an internal threat to the validity of our experiment. There are no well-defined standard guidelines for selecting the hyperparameters of the DRL algorithm. We followed the common practice of selecting hyperparameters to minimize the internal validity threat. We performed several experimental trials with different values of parameters and their combinations. At the end of these trials, we selected the best-suited hyperparameters for our experiment. To reduce the chances of construct validity threat, we used the same stopping criterion for our approach (AITester) and Random algorithm. That is, we execute AITester for 1000 episodes and Random for 1000 iterations during training. For the evaluation phase of both experiments, we repeated each approach 100 times which is a higher number as compared to traditional DRL evaluation procedure [25]. In the experiment with GCS-CDS, we configured and executed test paths according to the method defined in previous work [10].
@@ 5.6.3 Construct Validity Threat
Construct validity threat occurs when the relationship between cause and effect cannot be determined. To mitigate the possibility of this threat, we analyze the experimental results according to the guidelines provided by Arcuri and Briand [38]. We examine the experimental results based on their statistical significance and also look at the effect size. Specifically, we used the Wilcoxon signed-rank test for statistical significance and the Cliff’s Delta effect size measure [39] to analyze the effect size.
@@ 5.6.4 Conclusion Validity Threat
The threat to conclusion validity is associated with the effect of treatments on the outcomes. To reduce the chances of this threat, we use the moving average reward (MAR) metric to analyze the reward during the training session. Our approach utilizes UAV flight state machines to explore different flight behaviors. To analyze the diverse exploration of UAV flight states, we use path diversity as a metric that is used in a similar context and different domains, e.g., for Web [37]. We also used the unique number of OCL constraint violations as a metric to compare AITester with the CDST approach. For a fair comparison, this metric is common between this work and the previous work [10].
@@ 6.1 UAV Testing
Several approaches are available targeting different aspects such as tests generation from flight logs [54], analysis of boundary functions’ role in UAVs safety [55], faults identification using model checking [56], detecting errors in the physical unit of UAS [57], and mutation-based testing of robots [58]. Furthermore, some studies are conducted focusing on characterizing UAVs software faults such as [53, 59] and UAS autopilot faults [60]. Concerning these works, we focus on automating system-level testing of UAS.
@@ 6.2 Autopilot Testing
To test the autopilot of an aircraft, Fleck T. [61] manually designed various flight test scenarios considering the different ground and aerial situations such as the high speed at low-level, fly upwards, and altitude hold. Wells [62] presented various methods to test the flight control system of an autopilot. The proposed test methods target flight and missile stability control, and frequency response testing. Application and Flight Testing of an Adaptive Autopilot on Precision Guided Munitions Sharma et al. [63] manually developed nine flight tests to evaluate the performance of the proposed autopilot to support guided munitions. Kantue P. [64] manually created various flight test scenarios to test the autopilot system developed for missile control. To test Piccolo II autopilot, Anderson et al. [65] performed different flight tests using the manually designed test scenarios to check different conditions such as GPS range and radio communication. Ambroziak and Gosiewski [66] manually defined various flight tests to analyze the variation in airspeed, altitude, and trajectories followed by the UAV. Erdos and Watkins [67] manually created flight test scenarios to test the autopilot for the roll, pitch, altitude, airspeed, and navigation performance. Hartley et al. [68] manually developed seven flight test scenarios to analyze the waypoint navigation of the autopilot. In all the above-mentioned works, the test scenarios are created manually. Whereas our approach automatically performs system-level testing of the UAS. He et al. [69] proposed an approach to automatically generate test oracles for localizing faults in cyber-physical systems. The empirical evaluation using ArduPilot shows that the proposed approach generates better oracles as compared to humans. Afzal et al. [70] presented a technique to create oracles for cyber-physical systems based on historical telemetry data. The empirical evaluation using ArduPilot shows that the proposed technique outperforms the oracle generation approach by He et al. [69]. In comparison, our work focuses on the online testing of UAS using OCL-based test oracles. Recently, Leach et al. [71] presented a framework for assessing and handling the security situation during UAV operation. The evaluation using ArduPilot and Red Team attacks shows that the proposed framework can increase the dependability of autonomous systems. Our work targets the system-level testing of UAS instead of the security testing of UAVs.
@@ 6.4 Reinforcement Learning-based UAV Testing
A small amount of work has been done in the area of RL-based UAV testing. Junell et al. [75] devised an approach using reinforcement learning to reduce the number of trials during flight testing of a quadrotor. Siddiquee et al. [76] designed a vision-based reinforcement learning agent to test the guidance system of a quadcopter. Ma et al. [34] developed two algorithms based on reinforcement learning to test the self-healing part of cyber-physical systems under different uncertainties. The proposed algorithms were used to test the self-healing behavior of the ArduCopter. Ritchie Lee [77] presented Adastress, an approach for reinforcement learning-based stress testing of safety-critical systems. Subsequently, Moss et al. [78] extended Adastress for the flight trajectory planning system, incorporating the aspect of continuous decision-making into the reinforcement learning approach. While these works primarily focus on stress testing, our work mainly aims to automate system-level testing of UAS. Yuning He [79] introduced SYSAI, an approach that leverages AI techniques to automate various testing activities, such as property checking and test generation, for complex systems within the Aerospace domain. In comparison, our approach specifically targets UAS and employs reinforcement learning to automate system-level testing. In recent work, Zhang et al. [80] proposed a reinforcement learning-based approach for generating failure-causing inputs. The empirical evaluation results show the success of the proposed approach in failing the well-tested CPS. The main difference between these works with our work is that our approach applies DRL to automate the system-level testing of the UAV.
@@ 7 Limitations
In this section, we discuss the potential limitations of our study and outline prospective avenues for subsequent research.
@@ 7.4 Testing UAVs’ Swarm
The approach presented in this paper is a preliminary attempt to automate UAS testing. Thus, it is designed with a focus on a single subsystem of UAS as a SUT, such as a UAV or GCS at any given time. Our evaluation was also conducted considering testing UAV autopilot and GCS-CDS each as separate SUTs. Testing multiple UAS subsystems, such as a swarm of UAVs, is also a critical aspect of UAS testing. However, without empirical evidence, it remains uncertain whether our approach can be effectively applied to testing a swarm of UAVs or multiple GCSs. Further research is required to analyze the potential applicability of our approach to multiple UAS subsystems like UAVs’ swarm.
@@ 8 Conclusion
Automated system-level testing of the UAS is required to meet the testing guidelines set by different international safety standards. The prevailing industrial practice and the existing research works require manual effort in testing any of the UAS subsystems. Manual testing of a UAS subsystem is a tedious activity and only a handful of scenarios can be tested. In this paper, we propose a novel approach to automate the system-level testing of the UAS. The proposed approach utilizes MBT and AI techniques to create an AITester that generates, executes, and evaluates test scenarios based on the environmental context at runtime. We provide a toolset to support testing automation and to encourage further research and development. We perform two experiments to evaluate AITester using two different subsystems of the UAS. The first experiment is performed using the ArduCopter autopilot system and the second experiment is performed using an industrial case study of GCS-CDS. The results show that the AITester is effective in creating diverse test scenarios that lead to deviations from the expected behavior of the UAV autopilot and in revealing potential flaws in the GCS-CDS.
The approach presented in this paper is a fundamental step towards automating system-level testing of UAS. Our approach generates and executes test scenarios on the fly to find deviations from the expected behavior of UAS SUT. While the test scenarios obtained through this process can assist avionics testers in identifying faults during the initial development phases of UAS, they are not designed to qualify for test certifications specified by international safety standards. Furthermore, test certification qualifications are dependent on the practices of the avionics industry, specifically the safety standards they adhere to. In the future, we aim to concentrate on potential research directions, such as ensuring model correctness, learning state machines automatically, handling complex constraints, automating fault localization, and testing UAV swarms.
@@ 9 Acknowledgments
This research work is supported through a research grant titled ‘Establishment of National Center of Robotics and Automation (NCRA)’ by Higher Education Commission (HEC) Pakistan.

### v2

@@ 5.6.1 External Validity Threat
The possible threats to the validity of our experiment are discussed individually in the following subsections. The external threat to validity is related to the generalization of the experiment results. To reduce the chance of external validity threat, the experiment was performed using two primary subsystems of UAS, an autopilot (ArduCopter) and the CDS of GCS (GCS-CDS). ArduCopter is a good representative open-source autopilot system that is widely used in research and development [36, 17]. The ArduCopter supports a wide variety of multi-copter UAVs including a helicopter, tri-copter, quad-copter, hexacopter, etc. Moreover, this is also compatible with a wide range of hardware e.g., Pixhawk, CUAV v5 Plus, OpenPilot Revolution, etc. The GCS-CDS is an industrial case study containing widely used CDS elements. The results of our experiment may not be generalizable for all subsystems of UAS, however, this threat typically exists in every experiment [33].
@@ 5.6.2 Internal Validity Threat
The parameter tuning of the training algorithm presents an internal threat to the validity of our experiment. There are no well-defined standard guidelines for selecting the hyperparameters of the DRL algorithm. We followed the common practice of selecting hyperparameters to minimize the internal validity threat. We performed several experimental trials with different values of parameters and their combinations. At the end of these trials, we selected the best-suited hyperparameters for our experiment. To reduce the chances of construct validity threat, we used the same stopping criterion for our approach (AITester) and Random algorithm. We execute AITester for 1000 episodes and Random for 1000 iterations during training. For the evaluation phase of both experiments, we repeated each approach 100 times which is a higher number compared to traditional DRL evaluation procedure [28]. In the experiment with GCS-CDS, we configured and executed test paths according to the method defined in previous work [10].
@@ 5.6.3 Construct Validity Threat
Construct validity threat occurs when the relationship between cause and effect cannot be determined. To mitigate the possibility of this threat, we analyzed the experimental results according to the guidelines provided by Arcuri and Briand [40]. We examined the experimental results based on their statistical significance and also analyzed the effect size. Specifically, we used the Wilcoxon signed-rank test for statistical significance and the Cliff’s Delta effect size measure [41] to analyze the effect size.
@@ 5.6.4 Conclusion Validity Threat
The threat to conclusion validity is associated with the effect of treatments on the outcomes. To reduce the chances of this threat, we used the moving average reward (MAR) metric to analyze the reward during the training session. Our approach utilizes UAV flight state machines to explore different flight behaviors. To analyze the diverse exploration of UAV flight states, we used path diversity as a metric that was employed in a similar context and different domains, e.g., for Web [39]. We also used the unique number of OCL constraint violations as a metric to compare AITester with the CDST approach. For a fair comparison, this metric is common between this work and the previous work [10].
@@ 6.1 UAV Testing
Several approaches are available targeting different aspects such as tests generation from flight logs [56], analysis of boundary functions’ role in UAVs safety [57], faults identification using model checking [58], detecting errors in the physical unit of UAS [59], and mutation-based testing of robots [60]. Furthermore, some studies are conducted focusing on characterizing UAVs software faults such as [55, 61] and UAS autopilot faults [62]. In relation to these works, we focus on automating system-level testing of UAS.
@@ 6.2 Autopilot Testing
To test the autopilot of an aircraft, Fleck T. [63] manually designed various flight test scenarios considering the different ground and aerial situations such as the high speed at low-level, fly upwards, and altitude hold. Wells [64] presented various methods to test the flight control system of an autopilot. The proposed test methods target flight and missile stability control, and frequency response testing. Application and Flight Testing of an Adaptive Autopilot on Precision Guided Munitions Sharma et al. [65] manually developed nine flight tests to evaluate the performance of the proposed autopilot to support guided munitions. Kantue P. [66] manually created various flight test scenarios to test the autopilot system developed for missile control. To test Piccolo II autopilot, Anderson et al. [67] performed different flight tests using the manually designed test scenarios to check different conditions such as GPS range and radio communication. Ambroziak and Gosiewski [68] manually defined various flight tests to analyze the variation in airspeed, altitude, and trajectories followed by the UAV. Erdos and Watkins [69] manually created flight test scenarios to test the autopilot for the roll, pitch, altitude, airspeed, and navigation performance. Hartley et al. [70] manually developed seven flight test scenarios to analyze the waypoint navigation of the autopilot. In all the above-mentioned works, the test scenarios are created manually. Whereas our approach automatically performs system-level testing of the UAS. He et al. [71] proposed an approach to automatically generate test oracles for localizing faults in cyber-physical systems. The empirical evaluation using ArduPilot shows that the proposed approach generates better oracles compared to humans. Afzal et al. [72] presented a technique to create oracles for cyber-physical systems based on historical telemetry data. The empirical evaluation using ArduPilot shows that the proposed technique outperforms the oracle generation approach by He et al. [71]. In comparison, our work focuses on the online testing of UAS using OCL-based test oracles. Recently, Leach et al. [73] presented a framework for assessing and handling the security situation during UAV operation. The evaluation using ArduPilot and Red Team attacks shows that the proposed framework can increase the dependability of autonomous systems. Our work targets the system-level testing of UAS instead of the security testing of UAVs.
@@ 6.4 Reinforcement Learning-based UAV Testing
A small amount of work has been done in the area of RL-based UAV testing. Junell et al. [77] devised an approach using reinforcement learning to reduce the number of trials during flight testing of a quadrotor. Siddiquee et al. [78] designed a vision-based reinforcement learning agent to test the guidance system of a quadcopter. Ma et al. [36] developed two algorithms based on reinforcement learning to test the self-healing part of cyber-physical systems under different uncertainties. The proposed algorithms were used to test the self-healing behavior of the ArduCopter. Ritchie Lee [79] presented Adastress, an approach for reinforcement learning-based stress testing of safety-critical systems. Subsequently, Moss et al. [80] extended Adastress for the flight trajectory planning system, incorporating the aspect of continuous decision-making into the reinforcement learning approach. While these works primarily focus on stress testing, our work mainly aims to automate system-level testing of UAS. Yuning He [81] introduced SYSAI, an approach that leverages AI techniques to automate various testing activities, such as property checking and test generation, for complex systems within the Aerospace domain. In comparison, our approach specifically targets UAS and employs reinforcement learning to automate system-level testing. In recent work, Zhang et al. [82] proposed a reinforcement learning-based approach for generating failure-causing inputs. The empirical evaluation results show the success of the proposed approach in failing the well-tested CPS. The main difference between these works and our work is that our approach applies DRL to automate the system-level testing of the UAV.
@@ 7 Limitations
In this section, we discuss the potential limitations of our work and outline prospective avenues for subsequent research.
@@ 7.4 Testing UAVs’ Swarm
The approach presented in this paper is a preliminary attempt to automate UAS testing. Thus, it is designed with a focus on a single subsystem of UAS as a SUT, such as a UAV or GCS at any given time. Our evaluation was also conducted considering testing UAV autopilot and GCS-CDS each as separate SUTs. Testing multiple UAS subsystems, such as a swarm of UAVs, is also a critical aspect of UAS testing. However, without empirical evidence, it remains uncertain whether our approach can be effectively applied to testing a swarm of UAVs or multiple GCSs. Further research is required to analyze the potential applicability of our approach to multiple UAS subsystems like UAVs’ swarms.
@@ 8 Conclusion
Automated system-level testing of the UAS is required to meet the testing guidelines set by different international safety standards. The prevailing industrial practice and existing research works require manual effort in testing any of the UAS subsystems. Manual testing of a UAS subsystem is a tedious activity and only a handful of scenarios can be tested. In this paper, we proposed a novel approach to automate the system-level testing of the UAS. The proposed approach utilizes MBT and AI techniques to create an AITester that generates, executes, and evaluates test scenarios based on the environmental context at runtime. We provided a toolset to support testing automation and to encourage further research and development. We performed two experiments to evaluate AITester using two different subsystems of the UAS. The first experiment was performed using the ArduCopter autopilot system and the second experiment was performed using an industrial case study of GCS-CDS. The results demonstrated that the AITester is effective in creating diverse test scenarios that lead to deviations from the expected behavior of the UAV autopilot and in revealing potential flaws in the GCS-CDS.
The approach presented in this paper is a fundamental step towards automating system-level testing of UAS. Our approach generates and executes test scenarios on the fly to find deviations from the expected behavior of UAS SUT. While the test scenarios obtained through this process can assist avionics testers in identifying faults during the initial development phases of UAS, they are not designed to qualify for test certifications specified by international safety standards. Furthermore, test certification qualifications are dependent on the practices of the avionics industry, specifically the safety standards they adhere to.
In the future, we aim to concentrate on potential research directions, such as ensuring model correctness, learning state machines automatically, handling complex constraints, automating fault localization, and testing UAV swarms.
@@ 9 Acknowledgments
This research work is supported through a research grant titled ‘Establishment of National Center of Robotics and Automation (NCRA)’ by Higher Education Commission (HEC) Pakistan. We would like to acknowledge anonymous reviewers for their valuable suggestions, which greatly contributed to the significant improvement of this paper.